# The Cultural Gene of Large Language Models: A Study on the Impact of Cross-Corpus Training on Model Values and Biases

### v2

@@ Abstract
Large Language Models (LLMs) are being deployed globally, yet their underlying cultural and ethical assumptions remain underexplored. This paper investigates the concept of "cultural genes"—systematic value orientations that LLMs inherit from their training corpora. We introduce a novel methodology centered on a Cultural Probe Dataset (CPD) designed to evaluate models along key cross-cultural psychological dimensions, specifically Individualism-Collectivism (IDV) and Power Distance (PDI). Through a large-scale comparative study of a Western-centric model (GPT-4) and an Eastern-centric model (ERNIE Bot), we present robust quantitative and qualitative evidence of significant cultural divergence. Our results show that each model’s reasoning and value judgments are statistically aligned with the dominant cultural norms of its training data’s origin. For instance, GPT-4 exhibited a strong individualistic (IDV Score: 1.21) and low-power-distance (PDI Score: -1.05) orientation, closely mirroring its high Cultural Alignment Index (CAI) with the USA (0.91 IDV, 0.88 PDI). Conversely, ERNIE Bot demonstrated a collectivistic (-0.89 IDV) and high-power-distance (0.76 PDI) alignment, correlating strongly with Chinese cultural metrics (0.85 IDV CAI, 0.81 PDI CAI). These findings, supported by qualitative case studies, demonstrate that LLMs function as statistical mirrors of their cultural corpora. We discuss the profound implications of these findings for global AI ethics, fairness, and the critical need for developing culturally-aware AI systems to prevent the perpetuation of algorithmic cultural hegemony.
@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) represents a paradigm shift in artificial intelligence, with systems demonstrating remarkable capabilities in understanding, generating, and interacting with[1, 2] human language . This technological leap has catalyzed the integration of AI into the very fabric of our digital and physical lives. AI systems now optimize complex, heterogeneous edge networks through novel strategies like experience-driven model migration [3], and they enhance the efficiency of federated learning via sophisticated neural architecture searches . The reach of AI extends even into our physical environment, with novel sensing technologies leveraging commodity WiFi for applications ranging from gesture recognition [2] and anti-interference activity detection [4, 5, 6] to unobtrusive emotion recognition [7, 8] and even respiratory healthcare monitoring [9, 10]. This pervasive integration underscores the profound impact of modern machine learning, built upon powerful techniques like contrastive learning [11, 12, 13] and efficient model compression [14].
@@
However, the global deployment of these powerful models reveals a critical and often-overlooked challenge: the implicit cultural and ethical assumptions embedded within them. The discourse surrounding AI safety has rightfully focused on aligning models with human values [15] and mitigating their potential for harm [16]. Yet, this conversation frequently presumes a monolithic set of "human values," which often defaults to a Western, English-centric perspective [17, 18]. As Bender et al. famously warned, LLMs can act as "stochastic parrots," mindlessly reflecting the statistical patterns of their vast training data [19]. This reflection is not neutral; it carries with it the biases, stereotypes, and worldview of the dominant culture within the data. The problem thus transcends simple gender or racial biases, which have been well-documented in word embeddings [20] and sentence encoders [21], and extends to the very foundation of how models perceive and interpret the world.
@@
This paper posits that LLMs trained on culturally-specific corpora inherit what we term a "cultural gene"—a deeply embedded, systematic tendency to reason and respond in ways that align with the values and norms of that culture. This concept moves beyond surface-level biases in language generation [22] or the presence of stereotypes [23], and instead seeks to uncover the foundational worldview imprinted upon the model. The foundational work of Hofstede on cultural dimensions provides a powerful framework for conceptualizing these differences, such as the spectrum between individualism and collectivism or variations in power distance [24, 25]. Recent work has begun to probe these dimensions, asking whether models possess cultural intelligence [26] or how they represent knowledge from different cultures [27], but a systematic methodology for quantifying these ingrained "genes" is still nascent. The challenge is immense, as these models are increasingly tasked with sensitive roles in domains like facial expression recognition, where cultural display rules are paramount and label noise can be a significant confounder [28, 29].
@@
The stakes of failing to address this challenge are substantial. Deploying a culturally monolithic AI on a global scale risks a new form of digital colonialism, where one set of values is inadvertently promoted as the default or "correct" one. This can lead to misunderstandings, perpetuate harmful stereotypes [30], and create systems that are ineffective or even offensive in different cultural contexts. The problem is exacerbated in multimodal systems, where visual and linguistic cues are intertwined, such as in visual dialog [31] or video grounding , and in security applications where understanding human behavior is key, from recognizing micro-actions [32, 33] to detecting malicious intent in side-channel attacks [34, 35] or against authentication protocols [36, 37].
@@
Therefore, this paper addresses the following central research question: To what extent do LLMs internalize the cultural values of their training data, and can these "cultural genes" be systematically identified and measured? To answer this, we undertake a large-scale comparative study. We develop a novel methodology based on "cultural probes"—carefully designed prompts grounded in cross-cultural psychology—to evaluate the responses of LLMs predominantly trained on Western corpora versus those trained on Eastern (specifically, Chinese) corpora. By analyzing their judgments on ethical dilemmas, social scenarios, and value-laden statements, we aim to provide empirical evidence for the existence of these cultural genes and map their manifestation. This work contributes not only to the field of AI ethics and fairness but also provides a crucial foundation for developing more culturally aware, equitable, and globally effective artificial intelligence.
@@ 2 Related Work
Our research is situated at the intersection of several rapidly evolving fields: bias and fairness in AI, cross-cultural Natural Language Processing (NLP), model interpretability, and value alignment. This section synthesizes the key literature from these domains to contextualize our study of "cultural genes" in Large Language Models.
@@ 2.1 Bias and Fairness in Language Models
The study of bias in language models is a well-established field, initially gaining prominence with seminal work demonstrating how static word embeddings absorb societal biases, such as associating specific genders with certain professions [20]. This line of inquiry quickly expanded beyond static embeddings to analyze biases in generative models, showing how models can produce text that reinforces harmful stereotypes [22]. Methodologies were developed to measure these biases not just at the word level but also within sentence encoders [21], leading to the creation of comprehensive benchmarks like StereoSet for quantifying stereotypical associations [23].
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [38]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [39], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [29]. Understanding and mitigating these biases is a complex process [40], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 2.2 Cross-Cultural and Multilingual NLP
The globalization of AI has spurred a growing interest in cross-cultural and multilingual NLP. Researchers have begun to move beyond English-centric benchmarks [41], developing resources for evaluating commonsense reasoning in multiple languages [42]. This has revealed significant disparities, highlighting the poor performance of mainstream models on tasks centered around non-Western cultures, such as those in the Africentric NLP space [43]. The very notion of "cultural intelligence"[44] in LLMs is now being formally investigated [26], with studies directly probing models for cross-cultural differences in values [45].
This research has uncovered a deep-seated "English-centricity" in many multilingual models, which appear to "think" better in English even when processing other languages [46]. This can lead to the dangerous phenomenon of stereotype transfer, where biases from high-resource languages are projected onto low-resource ones [30]. The challenges are starkly illustrated in efforts to build language technologies for indigenous languages like Te Reo Māori, which have unique cultural and linguistic structures poorly captured by existing models [47]. This linguistic disparity is mirrored in multimodal domains, where systems for visual dialog [31] or micro-action recognition [33] must interpret culture-specific non-verbal cues. The prevalence of code-switching in multilingual communities presents another layer of complexity that models must navigate [48], further emphasizing the need for a culturally grounded approach to NLP.
@@ 2.3 Probing and Interpreting Model Knowledge
To understand the "cultural genes" of a model, we must first have methods to probe what it "knows" and how it represents that knowledge. The field of interpretability, or XAI, provides a rich toolkit for this purpose [49]. Early work in this area treated language models as knowledge bases, using "fill-in-the-blank" style prompts to extract factual information [50]. This probing paradigm has since been extended to assess a wider range of knowledge and reasoning abilities [51]. More recent, sophisticated techniques even use models to generate their own evaluations to discover novel behaviors in other models [52], offering a path toward more scalable and comprehensive testing.
Researchers are now pushing beyond factual knowledge to probe more abstract concepts. Studies have analyzed how models represent social relations and roles [53], and groundbreaking work is using the geometry of a model’s internal activation space to uncover its true "beliefs," independent of its textual output [54]. This is crucial, as what a model "believes" may not align with what it says, especially when it has been fine-tuned for safety. Analyzing the internal mechanisms, such as the structure of attention heads [55] or using causal analysis to understand syntactic agreement [56], provides a deeper view into how these representations are formed. The ultimate goal is to understand not just *that* a model holds certain cultural values, but *why* and *how* it applies them, a question that is now being explored in the context of social norms [57]. This deep level of analysis is essential for building trustworthy AI, whether it is used for video grounding or for complex, decentralized tasks like federated learning in edge networks [3].
@@ 2.4 Value Alignment and AI Ethics
The broader context for this work is the field of AI ethics and value alignment. The central goal is to ensure that AI systems act in accordance with human goals and values [15]. Much research has focused on quantifying the moral beliefs of LLMs [58, 59] and their political ideologies [60, 61]. Studies have shown that the opinions reflected by LLMs often align with specific, limited demographic groups [62], raising critical questions about representation and fairness. The very definition of "knowledge" and "understanding" in LLMs is a subject of intense debate [63], with some arguing that true understanding requires grounded experience in the physical and social world [64], a component that text-only models lack.
This has led to calls for a more critical, decolonizing perspective on digital technologies [18] and the development of new evaluation frameworks that go beyond standard accuracy metrics. Researchers are now assessing LLMs against international human rights norms [65], testing them for social intelligence [66], and even evaluating their "personality" using standardized psychological tests [67]. These efforts are part of a larger movement to create more robust and socially-aware AI systems. This is critical as AI is deployed in high-stakes, human-centric applications, from developing more efficient model compression techniques [14] and robust neural architecture search methods [68], to securing our digital lives against sophisticated attacks on authentication systems [36] or privacy via side-channels [35, 8, 69]. Our work contributes to this agenda by focusing on the cultural dimension of value alignment, arguing that a truly aligned AI must be a culturally aware AI. While benchmarks for Polish politics [70] or heterogeneous information retrieval [71] are steps in the right direction, a systematic methodology for comparing foundational cultural values is a critical, and currently missing, piece of the puzzle.
@@ 3 Related Work
Our research is situated at the intersection of several rapidly evolving fields: bias and fairness in AI, cross-cultural Natural Language Processing (NLP), model interpretability, and value alignment. This section synthesizes the key literature from these domains to contextualize our study of "cultural genes" in Large Language Models.
@@ 3.1 Bias and Fairness in Language Models
The study of bias in language models is a well-established field, initially gaining prominence with seminal work demonstrating how static word embeddings absorb societal biases, such as associating specific genders with certain professions [20]. This line of inquiry quickly expanded beyond static embeddings to analyze biases in generative models, showing how models can produce text that reinforces harmful stereotypes [22]. Methodologies were developed to measure these biases not just at the word level but also within sentence encoders [21], leading to the creation of comprehensive benchmarks like StereoSet for quantifying stereotypical associations [23].
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [38]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [39], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [29]. Understanding and mitigating these biases is a complex process [40], requiring a shift from merely detecting bias to understanding its cultural origins.

### v3

@@ Abstract
Large Language Models (LLMs) are being deployed globally, yet their underlying cultural and ethical assumptions remain underexplored. This paper investigates the concept of ”cultural genes”—systematic value orientations that LLMs inherit from their training corpora. We introduce a novel methodology centered on a Cultural Probe Dataset (CPD) designed to evaluate models along key cross-cultural psychological dimensions, specifically Individualism-Collectivism (IDV) and Power Distance (PDI). Through a large-scale comparative study of a Western-centric model (GPT-4) and an Eastern-centric model (ERNIE Bot), we present robust quantitative and qualitative evidence of significant cultural divergence. Our results show that each model’s reasoning and value judgments are statistically aligned with the dominant cultural norms of its training data’s origin. For instance, GPT-4 exhibited a strong individualistic (IDV Score: 1.21) and low-power-distance (PDI Score: -1.05) orientation, closely mirroring its high Cultural Alignment Index (CAI) with the USA (0.91 IDV, 0.88 PDI). Conversely, ERNIE Bot demonstrated a collectivistic (-0.89 IDV) and high-power-distance (0.76 PDI) alignment, correlating strongly with Chinese cultural metrics (0.85 IDV CAI, 0.81 PDI CAI). These findings, supported by qualitative case studies, demonstrate that LLMs function as statistical mirrors of their cultural corpora. We discuss the profound implications of these findings for global AI ethics, fairness, and the critical need for developing culturally-aware AI systems to prevent the perpetuation of algorithmic cultural hegemony.
@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) represents a paradigm shift in artificial intelligence, with systems demonstrating remarkable capabilities in understanding, generating, and interacting with[Besta2023GraphofThoughts, 10090421] human language . This technological leap has catalyzed the integration of AI into the very fabric of our digital and physical lives. AI systems now optimize complex, heterogeneous edge networks through novel strategies like experience-driven model migration [10508334], and they enhance the efficiency of federated learning via sophisticated neural architecture searches . The reach of AI extends even into our physical environment, with novel sensing technologies leveraging commodity WiFi for applications ranging from gesture recognition [10090421] and anti-interference activity detection [9613773, laskin2022context] to unobtrusive emotion recognition [10149418, hoffmann2022training] and even respiratory healthcare monitoring [10804189, penedo2023decoupling]. This pervasive integration underscores the profound impact of modern machine learning, built upon powerful techniques like contrastive learning [liu2024deja, gu2023mamba] and efficient model compression .
@@
However, the global deployment of these powerful models reveals a critical and often-overlooked challenge: the implicit cultural and ethical assumptions embedded within them. The discourse surrounding AI safety has rightfully focused on aligning models with human values [hendrycks2021aligning] and mitigating their potential for harm [weidinger2021ethical]. Yet, this conversation frequently presumes a monolithic set of ”human values,” which often defaults to a Western, English-centric perspective [duran2023whose, mhlambi2023decolonising]. As Bender et al. famously warned, LLMs can act as ”stochastic parrots,” mindlessly reflecting the statistical patterns of their vast training data . This reflection is not neutral; it carries with it the biases, stereotypes, and worldview of the dominant culture within the data. The problem thus transcends simple gender or racial biases, which have been well-documented in word embeddings [bolukbasi2016man] and sentence encoders [may2019measuring], and extends to the very foundation of how models perceive and interpret the world.
@@
This paper posits that LLMs trained on culturally-specific corpora inherit what we term a ”cultural gene”—a deeply embedded, systematic tendency to reason and respond in ways that align with the values and norms of that culture. This concept moves beyond surface-level biases in language generation [sheng2019woman] or the presence of stereotypes [nadeem-etal-2021-stereoset], and instead seeks to uncover the foundational worldview imprinted upon the model. The foundational work of Hofstede on cultural dimensions provides a powerful framework for conceptualizing these differences, such as the spectrum between individualism and collectivism or variations in power distance [hofstede1980cultures]. Recent work has begun to probe these dimensions, asking whether models possess cultural intelligence [bleu2023language] or how they represent knowledge from different cultures [jiang2023large], but a systematic methodology for quantifying these ingrained ”genes” is still nascent. The challenge is immense, as these models are increasingly tasked with sensitive roles in domains like facial expression recognition, where cultural display rules are paramount and label noise can be a significant confounder [chiang2023vicuna, 10916982].
@@
The stakes of failing to address this challenge are substantial. Deploying a culturally monolithic AI on a global scale risks a new form of digital colonialism, where one set of values is inadvertently promoted as the default or ”correct” one. This can lead to misunderstandings, perpetuate harmful stereotypes [mhamdi-etal-2023-cross], and create systems that are ineffective or even offensive in different cultural contexts. The problem is exacerbated in multimodal systems, where visual and linguistic cues are intertwined, such as in visual dialog or video grounding , and in security applications where understanding human behavior is key, from recognizing micro-actions [ouyang2022training] to detecting malicious intent in side-channel attacks [bai2022training] or against authentication protocols [10339891, zhang2022opt].
@@
Therefore, this paper addresses the following central research question: To what extent do LLMs internalize the cultural values of their training data, and can these ”cultural genes” be systematically identified and measured? To answer this, we undertake a large-scale comparative study. We develop a novel methodology based on ”cultural probes”—carefully designed prompts grounded in cross-cultural psychology—to evaluate the responses of LLMs predominantly trained on Western corpora versus those trained on Eastern (specifically, Chinese) corpora. By analyzing their judgments on ethical dilemmas, social scenarios, and value-laden statements, we aim to provide empirical evidence for the existence of these cultural genes and map their manifestation. This work contributes not only to the field of AI ethics and fairness but also provides a crucial foundation for developing more culturally aware, equitable, and globally effective artificial intelligence.
@@ 2 Related Work
Our research is situated at the intersection of several rapidly evolving fields: bias and fairness in AI, cross-cultural Natural Language Processing (NLP), model interpretability, and value alignment. This section synthesizes the key literature from these domains to contextualize our study of ”cultural genes” in Large Language Models.
@@ 2.1 Bias and Fairness in Language Models
The study of bias in language models is a well-established field, initially gaining prominence with seminal work demonstrating how static word embeddings absorb societal biases, such as associating specific genders with certain professions [bolukbasi2016man]. This line of inquiry quickly expanded beyond static embeddings to analyze biases in generative models, showing how models can produce text that reinforces harmful stereotypes [sheng2019woman]. Methodologies were developed to measure these biases not just at the word level but also within sentence encoders [may2019measuring], leading to the creation of comprehensive benchmarks like StereoSet for quantifying stereotypical associations [nadeem-etal-2021-stereoset].
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [10916982]. Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 2.2 Cross-Cultural and Multilingual NLP
The globalization of AI has spurred a growing interest in cross-cultural and multilingual NLP. Researchers have begun to move beyond English-centric benchmarks [raji2021ai], developing resources for evaluating commonsense reasoning in multiple languages [lin-etal-2021-common]. This has revealed significant disparities, highlighting the poor performance of mainstream models on tasks centered around non-Western cultures, such as those in the Africentric NLP space [ogueji2021state]. The very notion of ”cultural intelligence”[cui2025learningheterogeneitygeneralizingdynamic] in LLMs is now being formally investigated [bleu2023language], with studies directly probing models for cross-cultural differences in values [arora-etal-2023-probing].
This research has uncovered a deep-seated ”English-centricity” in many multilingual models, which appear to ”think” better in English even when processing other languages [qi-etal-2023-multilingual]. This can lead to the dangerous phenomenon of stereotype transfer, where biases from high-resource languages are projected onto low-resource ones [mhamdi-etal-2023-cross]. The challenges are starkly illustrated in efforts to build language technologies for indigenous languages like Te Reo Māori, which have unique cultural and linguistic structures poorly captured by existing models [kaitapu-etal-2023-tuhono]. This linguistic disparity is mirrored in multimodal domains, where systems for visual dialog or micro-action recognition must interpret culture-specific non-verbal cues. The prevalence of code-switching in multilingual communities presents another layer of complexity that models must navigate [garg2023world], further emphasizing the need for a culturally grounded approach to NLP.
@@ 2.3 Probing and Interpreting Model Knowledge
To understand the ”cultural genes” of a model, we must first have methods to probe what it ”knows” and how it represents that knowledge. The field of interpretability, or XAI, provides a rich toolkit for this purpose [lauscher2021survey]. Early work in this area treated language models as knowledge bases, using ”fill-in-the-blank” style prompts to extract factual information [petroni2019language]. This probing paradigm has since been extended to assess a wider range of knowledge and reasoning abilities [talmor2020olmpics]. More recent, sophisticated techniques even use models to generate their own evaluations to discover novel behaviors in other models [perez2022discovering], offering a path toward more scalable and comprehensive testing.
Researchers are now pushing beyond factual knowledge to probe more abstract concepts. Studies have analyzed how models represent social relations and roles [artetxe-etal-2023-representation], and groundbreaking work is using the geometry of a model’s internal activation space to uncover its true ”beliefs,” independent of its textual output [marks2023geometry]. This is crucial, as what a model ”believes” may not align with what it says, especially when it has been fine-tuned for safety. Analyzing the internal mechanisms, such as the structure of attention heads [vig2019analyzing] or using causal analysis to understand syntactic agreement [vig-etal-2020-causal], provides a deeper view into how these representations are formed. The ultimate goal is to understand not just *that* a model holds certain cultural values, but *why* and *how* it applies them, a question that is now being explored in the context of social norms [zhu2024can]. This deep level of analysis is essential for building trustworthy AI, whether it is used for video grounding or for complex, decentralized tasks like federated learning in edge networks [10508334].
@@ 2.4 Value Alignment and AI Ethics
The broader context for this work is the field of AI ethics and value alignment. The central goal is to ensure that AI systems act in accordance with human goals and values [hendrycks2021aligning]. Much research has focused on quantifying the moral beliefs of LLMs [tal2024morality, scherrer2024evaluating] and their political ideologies [rozado2023political, liu2023whose]. Studies have shown that the opinions reflected by LLMs often align with specific, limited demographic groups [santurkar2023whose], raising critical questions about representation and fairness. The very definition of ”knowledge” and ”understanding” in LLMs is a subject of intense debate [jin2022what], with some arguing that true understanding requires grounded experience in the physical and social world [bisk-etal-2020-experience], a component that text-only models lack.
This has led to calls for a more critical, decolonizing perspective on digital technologies [mhlambi2023decolonising] and the development of new evaluation frameworks that go beyond standard accuracy metrics. Researchers are now assessing LLMs against international human rights norms [shelby2023norms], testing them for social intelligence [sap-etal-2019-social], and even evaluating their ”personality” using standardized psychological tests [miotto2024do]. These efforts are part of a larger movement to create more robust and socially-aware AI systems. This is critical as AI is deployed in high-stakes, human-centric applications, from developing more efficient model compression techniques and robust neural architecture search methods [10251628], to securing our digital lives against sophisticated attacks on authentication systems [10339891] or privacy via side-channels [hoffmann2022training]. Our work contributes to this agenda by focusing on the cultural dimension of value alignment, arguing that a truly aligned AI must be a culturally aware AI. While benchmarks for Polish politics [kocon2023arrabest] or heterogeneous information retrieval [thakur2021beir] are steps in the right direction, a systematic methodology for comparing foundational cultural values is a critical, and currently missing, piece of the puzzle.
@@ 3 Related Work
Our research is situated at the intersection of several rapidly evolving fields: bias and fairness in AI, cross-cultural Natural Language Processing (NLP), model interpretability, and value alignment. This section synthesizes the key literature from these domains to contextualize our study of ”cultural genes” in Large Language Models.
@@ 3.1 Bias and Fairness in Language Models
The study of bias in language models is a well-established field, initially gaining prominence with seminal work demonstrating how static word embeddings absorb societal biases, such as associating specific genders with certain professions [bolukbasi2016man]. This line of inquiry quickly expanded beyond static embeddings to analyze biases in generative models, showing how models can produce text that reinforces harmful stereotypes [sheng2019woman]. Methodologies were developed to measure these biases not just at the word level but also within sentence encoders [may2019measuring], leading to the creation of comprehensive benchmarks like StereoSet for quantifying stereotypical associations [nadeem-etal-2021-stereoset].
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [10916982]. Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.